Excel Online Structured Activity: CAPM, portfolio risk, and return Consider the following information for three stocks, StockHi! I need help with A, B, and C, please. Thanks :)
Answer #1

Please refer to below spreadsheet for calculation and answer. Cell reference also provided.

A B C D E Stock LA Beta 0.8 Expected Return 8.78% 10.01% 11.65% Standard deviation 15% 15% 15% Weight in Fund P 33.33% 33.33%

Cell reference -

B с Expected Return Stock A B C 0.0878 0.1001 0.1165 Standard deviation Beta Weight in Fund P 0.15 0.8 =1/3 0.15 1.1 =1/3 0.1

Hope this will help, please do comment if you need any further explanation. Your feedback would be highly appreciated.